OpenStudy (anonymous):

Please Help !!!! Jessie and Bob are financing $425,500 to purchase a house. They obtained a 30/8 balloon mortgage at 6.55%. What will their balloon payment be? $379,659.89 $637,130.99 $357,491.39 $380,040.19

OpenStudy (amistre64):

calculate the monthly payment 1000k^(12*30)(1-k)/(1-k^(12*30)), k=1+.0655/12 payments are 6.3536 per 1000 loaned http://www.wolframalpha.com/input/?i=1000k%5E%2812*30%29%281-k%29%2F%281-k%5E%2812*30%29%29%2C+k%3D1%2B.0655%2F12

OpenStudy (amistre64):

using that we can determine the balance owed after 8 years B8 = B (k^(12*8) -(6.3536)(1-k^(12*8))/(1-k), k=1+.0655/12 http://www.wolframalpha.com/input/?i=425500k%5E%2812*8%29+-+2703.46%281-k%5E%2812*8%29%29%2F%281-k%29%2C+k%3D1%2B.0655%2F12

OpenStudy (anonymous):

im still confused